How do you solve the partial differential equation (3D^² +10DD^1+3D ^1^2) z = e^x-y?


1
Expert's answer
2022-02-08T15:27:29-0500

we shall solve the partial differential equation

"(3D^{2}+10DD^{\prime}+3D^{\prime^{2}})z=e^{x-y}"

The auxilliary equation is

"3m^{2}+10m+3=0\\(3m+1)(m+3)=0"

"3m+1=0" or "m+3=0"

"m=-\frac{1}{3},3"


"C.F=f_{1}(y-\frac{1}{3}x)+f_{2}(y-3x)"


We have the P.I to be

"P.I=\frac{1}{3D^{2}+10DD^{\prime}+3D^{\prime^{2}}}e^{x-y}"


"=\frac{1}{3(1)^{2}+10(1)(-1)+3(-1)^{2}}e^{x-y}"


"=\frac{1}{3-10+3}e^{x-y}"


"=-\frac{1}{4}e^{x-y}"

"z=f_{1}(y-\frac{1}{3}x)+f_{2}(y-3x)-\frac{1}{4}e^{x-y}"

which is the general solution for the differential equation.